OpenStudy (anonymous):

Choose the correct simplification of (7x + 4)(9x – 3)

OpenStudy (imtant):

Expand bothe the brackets 63x^2-21x+36x-12=0 63x^2+15x-12=0 So,it is (3x-1)(7x+4)=0 The values of x are 1/3 and -4/7.

OpenStudy (anonymous):

so its 63x2 + 15x + 12?

OpenStudy (imtant):

yes it is.if the question say 'solve', you need to find out the values of x.

OpenStudy (anonymous):

so we add the 12 not minus

OpenStudy (imtant):

Sorry.t is minus not +.4*-3 is -12.

OpenStudy (imtant):

*it is minus